You are here: Foswiki>PartnerObserving Web>ItalyQueue2013>ObsInaf20130926 (29 Sep 2013, EleanoraSani)Edit Attach

INAF queue observing, 2013 Sep 25-26

Observer: E. Sani, F. Cusano
Telescope Operator: D. Gonzales-Huerta
Support Astronomer: D. Thompson
Instrument: LUCI

Summary

MODS1 is currently not available due to problems with the red CCD electronics.
At sunset we have to remain closed due to strong wind.
At the beginning of the observations the sky is clear.
At 5.00 UT we need to close the dome due to strong wind. When the TO tries to open again, the left door turns to be stuck. We need to call J. Urban.
The left door problem was solved, but wind gusts increased again. The TO closed the dome.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Cresci_8B DSF2237b K, 210_zJHK 20 min within the specs
Cresci_8B Hip104320 K, 210_zJHK    

Details

03:16 the TO opens the dome

03:50 slewing to telluric standard of Cresci
executeLUCI tell2B_mos.acq luci.20130926.0041-0043 (we are using the N3.75 camera for the acquisition)

04:00 executeLUCI tellB_mos_210_K1.spec 0044-0047 (too bright, move to another standard)

04:07 executeLUCI tellB_mos.acq 0048-0049

04:12 executeLUCI tellB_mos_210_K1.spec 0050-0053

04:16 slewing to the target Cresci_B (seeing 0.85") executeLUCI campoB_mos.acq 0054-0057

04:32 executeLUCI campoB_mos_210_K_4.spec 0058-0061
  • average seeing is 0.8 on the GCS

04:55 the TO closes the dome because of the strong wind

07.40 the TO opens the dome
  • the left door is stuck, we need to call J. Urban

08.35 We tried to acquire the standard but there were again strong wind gusts. The TO close the dome

-- EleanoraSani - 26 Sep 2013
Topic revision: r6 - 29 Sep 2013, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback